### v1

@@ 1 Introduction
Stochastic approximation (SA), introduced by (Robbins and Monro, 1951), is an iterative sample-based approach to find the root (or fixed point) $x^{\star}$ of some unknown operator $F$. In particular, SA iteratively updates $x$, an estimate of $x^{\star}$, by moving along the direction of the sample $F(x;\xi)$ scaled by some step size, where $\xi$ is a random variable representing sampling noise. The simplicity of its implementation has enabled broad applications in many areas including stochastic optimization, machine learning, and reinforcement learning (Sutton and Barto, 2018; Lan, 2020). The convergence properties of SA are well-studied, where the error rate achieved by SA with Polyak-Ruppert (PR) averaging is known to decay faster than that of SA without PR averaging (Polyak and Juditsky, 1992).
@@
Unlike the literature on SA where convergence properties are well understood, theoretical results for TSA are not complete. While the benefit of using PR averaging has been established for SA (Polyak and Juditsky, 1992), its benefit has only been demonstrated with respect to asymptotic convergence in distribution for TSA (Mokkadem and Pelletier, 2006). In this paper, we focus on understanding the finite-time error rates at which the PR average of the iterates generated by TSA converge to the desired solution. We use TSA to refer to the sequence $\{(x_{t},y_{t})\}$ in Eq. (4) without Polyak-Ruppert averaging, to distinguish from TSA-PR which is used to refer to the averages $(\bar{x}_{n},\bar{y}_{n})$. Our contributions are summarized below.
@@ 2 Model and Preliminaries
$A_{ff}+A_{ff}^{T}\succ 0,\quad\Delta+\Delta^{T}\succ 0.$ (8)
When a matrix $-A$ is Hürwitz stable, we use $\mu_{A},\nu_{A}$ to denote the smallest and largest eigenvalues of $A+A^{T}$, respectively.
@@
Here we allow arbitrary $\alpha_{1},\gamma_{1}>0$, but clarify a few technical conditions. Define $(\mu_{ff},\nu_{ff})$ and $(\mu_{\Delta},\nu_{\Delta})$ such that $\mu_{ff}I\preceq A_{ff}+A_{ff}^{T}\preceq\nu_{ff}I$ and $\mu_{\Delta}I\preceq\Delta+\Delta^{T}\preceq\nu_{\Delta}I$. Because $\alpha_{t},\gamma_{t}\to 0$ and $\gamma_{t}/\alpha_{t}\to 0$, there exists a time $t_{0}$ and problem-dependent constants $K_{1},K_{2},K_{3}$ such that the following holds for all $t\geq t_{0}$:
@@ 3 Main Results
Recall the notations $\hat{x}_{n}=x_{n}-x^{*},\hat{y}_{n}=y_{n}-y^{*}$ to denote the deviation of TSA $(x_{n},y_{n})$ from the solution $(x^{*},y^{*})$.
@@
$\begin{split}\lVert\mathbb{E}\hat{x}_{n+1}\hat{x}_{n+1}^{T}-\alpha_{n+1}\Sigma_{ff}\rVert&\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert\mathbb{E}\hat{x}_{1}\hat{x}_{1}^{T}-\alpha_{1}\Sigma_{ff}\rVert+M_{f}\gamma_{n},\\ \lVert\mathbb{E}\hat{y}_{n+1}\hat{y}_{n+1}^{T}-\gamma_{n+1}\Sigma_{ss}\rVert&\leq\prod_{t=1}^{n}\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert\mathbb{E}\hat{y}_{1}\hat{y}_{1}^{T}-\gamma_{1}\Sigma_{ss}\rVert+\frac{M_{s}}{n}\end{split}$ (11)
@@
The limiting distribution of TSA-PR involves the Schur complement $G=A_{ff}-A_{fs}A_{ss}^{-1}A_{ss}$ of $A_{ss}$, as well as the Schur complement $\Delta$ defined in Section 2.
@@
$\begin{split}d_{1}\left(\sqrt{n}G(\bar{x}_{n}-x^{*}),(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(n^{a/2}+n^{a-b/2}+n^{b/2}\right)\right),\\ d_{1}\left(\sqrt{n}\Delta(\bar{y}_{n}-y^{*}),(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(n^{a/2}+n^{a-b/2}+n^{b/2}\right)\right),\end{split}$ (15)
@@
A closed-form solution for $\bar{\Sigma}_{ff}$ and $\bar{\Sigma}_{ss}$ is given in Section C, Eq. (61). Because $a<b$, the dominant term is $n^{b/2}$, but we explicitly state the dependence on $a,b$, and $a-b/2$ which decays with $\gamma_{t}/\alpha_{t}$ to illustrate their roles. By restricting the test functions $h$ in the Wasserstein-1 distance in Eq. (13) to be $h(x)\in\{-\lVert x\rVert,\lVert x\rVert\}$, we have the following lower and upper bounds on the expected error.
@@
The bounds on the expected error is tight, in the sense that the expected error is bounded above by rate $n^{-1/2}$, and similar to Eq. (12), both $\mathbb{E}\lVert\bar{x}_{n}-x^{*}\rVert$ and $\mathbb{E}\lVert\bar{y}_{n}-y^{*}\rVert$ are lower bounded by the expected norm of the limiting Gaussian vector times the rate $n^{-1/2}$. Replacing the norm of the Gaussian vector with the trace of its covariance would yield $\mathcal{O}(n^{-1/2})$ instead of the $o(n^{-1/2})$ above. Theorem 3.2 shows that the step sizes determine the quality of a Gaussian approximation of $\bar{z}_{n}-z^{*}$ at any $n\geq 1$, whereas Corollary 3.3 shows that the expected error decays at rate $n^{-1/2}$, with the step sizes $11/2<a<b<$ 1 determining only the transient behavior.
@@ 4 Experiments
The expected errors of $x_{t}$ and $y_{t}$ for every $t\leq n$ are recovered from Theorem 3.1 and Jensen’s inequality, which gives $\mathbb{E}\lVert x_{t}\rVert=\mathcal{O}(t^{-a/2})$ and $\mathbb{E}\lVert y_{t}\rVert=\mathcal{O}(t^{-b/2})$. Combining, we conclude that the sum of telescoped sums in Eq. (34) grows with $n^{a/2},n^{b/2}$, and $n^{a-b/2}$, where the $a-b/2$ is obtained from the $\gamma_{n+1}$ dependence of $\mathbb{E}\lVert\hat{x}_{n}\rVert^{2}$ in Eq. (31).
@@
$\begin{split}\mathbb{E}\lVert G\bar{x}_{n}\rVert-n^{-1/2}\mathbb{E}\lVert(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\rVert&\leq d_{1}\left(G\bar{x}_{n},n^{-1/2}(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right)=\mathcal{O}(n^{b/2-1}),\\ \mathbb{E}\lVert\Delta\bar{y}_{n}\rVert-n^{-1/2}\mathbb{E}\lVert(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\rVert&\leq d_{1}\left(\Delta\bar{y}_{n},n^{-1/2}(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\right)=\mathcal{O}(n^{b/2-1}).\end{split}$ (35)
Because $b<1$, we have that $\mathbb{E}\lVert G\bar{x}_{n}\rVert$ and $\mathbb{E}\lVert\Delta\bar{y}_{n}\rVert$ are both $\mathcal{O}(n^{-1/2})$. The lower bound is obtained by using the test function $-\lVert x\rVert$.
@@ A.1 Recursion Updates
Second, we design a sequence of matrices $\{L_{t}\}$ such that the recursion for $\tilde{x}_{t}=\hat{x}_{t}+L_{t}\hat{y}_{t}$ does not depend on $\hat{y}_{t}$ explicitly. The choice of $L_{t}$ determines the coordinate transformation
@@ A.2 Usage of Assumptions 13
Here we state a few elementary properties that will be used to prove Theorems 3.1 and 3.2. We will express the difference between finite-time second moments and the asymptotic covariance as a contraction, and so we require that the contraction is well defined for symmetric matrices, rather than for definite matrices. The contraction that appears often is as follows. Given any matrix $A$, we denote $\|A\|$ is the matrix norm of $A$. Let $A$ satisfy $\mu I\preceq A+A^{T}\preceq\nu I$. Given any symmetric matrix $R$, we have
$R-\alpha(AR+RA^{T})$ $=(I-\alpha A)R(I-\alpha A)^{T}-\alpha^{2}ARA^{T},$
which implies
$\lVert R-\alpha(AR+RA^{T})\rVert$ $\leq\lVert I-\alpha A\rVert^{2}\lVert R\rVert+\alpha^{2}\lVert ARA^{T}\rVert.$
In addition, we have $\lVert A\rVert\leq\nu/2$, which gives
$\lVert I-\alpha A\rVert^{2}=\sup_{x\neq 0}\left(1-\alpha\frac{x^{T}(A+A^{T})x}{\lVert x\rVert^{2}}+\alpha^{2}\frac{\lVert Ax\rVert^{2}}{\lVert x\rVert^{2}}\right)\leq 1-\alpha\mu+\alpha^{2}\lVert A\rVert^{2}.$
Thus, for $\alpha\leq\mu/\nu^{2}$ we obtain
$\lVert R-\alpha(AR+RA^{T})\rVert\leq\left(1-\alpha\mu+\alpha^{2}\frac{\nu^{2}}{2}\right)\lVert R\rVert\leq\left(1-\alpha\frac{\mu}{2}\right)\lVert R\rVert.$ (41)
@@ A.3 Size of Linear Transformation
It was shown in Lemma 18 of (Kaledin et al., 2020) that $I-\gamma_{t}B_{t}^{ss}$ is non-singular and that $\lVert L_{t}\rVert$ is uniformly bounded. Therefore, $\max_{t\leq n}\prod_{j=t+1}^{n}\lVert(I-\gamma_{t}B_{t}^{ss})^{-1}\rVert$ is uniformly bounded by some constant $K_{ss}$ and we have
$\lVert L_{t+1}\rVert$ $\leq\lVert I-\alpha_{t}A_{ff}\rVert\lVert L_{t}(I-\gamma_{t}B_{t}^{ss})^{-1}\rVert+\gamma_{t}\lVert A_{ff}^{-1}A_{fs}B_{t}^{ss}\rVert$
$\leq\sqrt{1-\alpha_{t}\frac{\mu_{ff}}{2}}\lVert(I-\gamma_{t}B_{n}^{ss})^{-1}\rVert\lVert L_{t}\rVert+\gamma_{t}\lVert A_{ff}^{-1}A_{fs}B_{n}^{ss}\rVert$
$\leq\left(1-\frac{\mu_{ff}}{4}\alpha_{t}\right)\lVert(I-\gamma_{t}B_{t}^{ss})^{-1}\rVert\lVert L_{t}\rVert+\gamma_{t}\lVert A_{ff}^{-1}A_{fs}B_{t}^{ss}\rVert$
when $\alpha_{t}\leq\mu/\nu^{2}$, where we used that $\sqrt{1-x}\leq 1-x/2$ when $x\in[0,1]$. By induction and Eq. (42), there exists a constant $K_{L}$ such that
$\frac{1}{K_{ss}}\lVert L_{n+1}\rVert$ $\leq\prod_{t=1}^{n}\left(1-\frac{\mu_{ff}}{4}\alpha_{t}\right)\lVert L_{1}\rVert+\sum_{t=1}^{n}\gamma_{t}\prod_{j=t+1}^{n}\left(1-\frac{\mu_{ff}}{4}\alpha_{j}\right)\lVert A_{ff}^{-1}A_{fs}B_{t}^{ss}\rVert$
$\leq\prod_{t=1}^{n}\left(1-\frac{\mu_{ff}}{4}\alpha_{t}\right)\lVert L_{1}\rVert+K_{L}\frac{\gamma_{n}}{\alpha_{n}}.$
Appendix B Mean Square Analysis of TTSA Last Iterates
@@ B.1 Fast-Time-Scale Mean-Square Convergence
We note that $\Sigma_{ff}$ exists since $A_{ff}$ satisfies $A_{ff}+A_{ff}^{T}$ is negative.
@@
$\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert\mathbb{E}X_{1}-\alpha_{1}\Sigma_{ff}\rVert+M_{f}\gamma_{n}.$
@@
$\lVert R_{t}^{ff}(X_{t})\rVert\leq M_{f}^{\prime}(\lVert X_{t}\rVert+\alpha_{t})\leq M_{f}^{\prime}\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert+M_{f}^{\prime}\alpha_{t}(1+\lVert\Sigma_{ff}\rVert).$
Such a constant exists because $X_{t}$ is positive definite and $\lVert\delta_{t}\rVert$ in (46) is uniformly bounded by Proposition A.1. The choice of step sizes ensures Eq. (41) is satisfied for the symmetric matrix $X_{t}-\alpha_{t}\Sigma_{ff}$. By the property $\alpha_{t}-\alpha_{t+1}\leq t^{-1}\alpha_{t+1}$,
$\lVert X_{t+1}-\alpha_{t+1}\Sigma_{ff}\rVert$ $\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{2}+\gamma_{t}M_{f}^{\prime}\right)\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert+\frac{\alpha_{t+1}}{t}\lVert\Sigma_{ff}\rVert$
$+\alpha_{t}\gamma_{t}M_{f}^{\prime}(\lVert\Sigma_{ff}\rVert+1)$
$\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert\tilde{x}_{t}\tilde{x}_{t}^{T}-\alpha_{t}\Sigma_{ff}\rVert+\frac{\alpha_{t}}{t}\lVert\Sigma_{ff}\rVert+\alpha_{t}\gamma_{t}M_{f}^{\prime}(\lVert\Sigma_{ff}\rVert+1).$
The above equation is of the form
$\lVert X_{t+1}-\alpha_{t+1}\Sigma_{ff}\rVert\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert+M_{f}^{\prime\prime}\alpha_{t}\gamma_{t},$ (50)
where $M_{f}^{\prime\prime}$ subsumes the coefficients. By induction and Eq. (42), we have that for some constant $M_{f}>0$, it holds that
$\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert$ $\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{1}-\alpha_{1}\Sigma_{ff}\rVert+M_{f}^{\prime\prime}\sum_{t=1}^{n}\alpha_{t}\gamma_{t}\prod_{j=t+1}^{n}\left(1-\alpha_{j}\frac{\mu_{ff}}{4}\right)$
$\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{1}-\alpha_{1}\Sigma_{ff}\rVert+M_{f}\gamma_{n}.$

### v2

@@ 1 Introduction
Stochastic approximation (SA), introduced by Robbins and Monro (1951), is an iterative sample-based approach to find the root (or fixed point) $x^{\star}$ of some unknown operator $F$. In particular, SA iteratively updates $x$, an estimate of $x^{\star}$, by moving along the direction of the sample $F(x;\xi)$ scaled by some step size, where $\xi$ is a random variable representing sampling noise. The simplicity of its implementation has enabled broad applications in many areas including stochastic optimization, machine learning, and reinforcement learning (Sutton and Barto, 2018; Lan, 2020). The convergence properties of SA are well-studied, where the error rate achieved by SA with Polyak-Ruppert (PR) averaging is known to decay faster than that of SA without PR averaging (Polyak and Juditsky, 1992).
@@
Unlike the literature on SA where convergence properties are well understood, theoretical results for TSA are not complete. While the benefit of using PR averaging has been established for SA (Polyak and Juditsky, 1992), its benefit has only been demonstrated with respect to asymptotic convergence in distribution for TSA (Mokkadem and Pelletier, 2006). In this paper, we focus on understanding the finite-time error rates at which the PR average of the iterates generated by TSA converge to the desired solution. We use TSA to refer to the sequence $\{(x_{t},y_{t})\}$ in Eq. (4) without Polyak-Ruppert averaging, to distinguish from TSA-PR which is used to refer to the averages $\bar{x}_{n}=n^{-1}\sum_{t=1}^{n}x_{t}$ and $\bar{y}_{n}=n^{-1}\sum_{t=1}^{n}y_{t}$. Our contributions are summarized below.
@@ 2 Model and Preliminaries
$A_{ff}P_{ff}+P_{ff}A_{ff}^{T}=I,\quad\Delta P_{\Delta}+P_{\Delta}\Delta^{T}=I$ (8)
for some positive definite matrices $P_{ff}$ and $P_{\Delta}$. The operator norms of $P_{ff}$ and $P_{\Delta}$ are denoted as $\nu_{ff}$ and $\nu_{\Delta}$, respectively.
The above condition implies a contraction in the weighted norm
$\lVert I-2\alpha A_{ff}\rVert_{P_{ff}}\leq\left(1-\frac{\mu_{ff}}{2}\alpha\right)$ (9)
for some positive constant $\mu_{ff}$ and any sufficiently small $\alpha>0$. The constants $\mu_{ff},\mu_{\Delta}$ are described in more detail in Section A.2.
@@
Here we allow arbitrary $\alpha_{1},\gamma_{1}>0$, but clarify a few technical conditions. Recall the constants $(\mu_{ff},\nu_{ff})$ and $(\mu_{\Delta},\nu_{\Delta})$ associated with $A_{ff}$ and $\Delta$ from Assumption 1. Because $\alpha_{t},\gamma_{t}\to 0$ and $\gamma_{t}/\alpha_{t}\to 0$, there exists a time $t_{0}$ and problem-dependent constants $K_{1},K_{2},K_{3}$ such that the following holds for all $t\geq t_{0}$:
@@ 3 Main Results
Recall the notations $\hat{x}_{n}=x_{n}-x_{\infty}(y_{n}),\hat{y}_{n}=y_{n}-y^{*}$ to denote the deviation of TSA $(x_{n},y_{n})$ from their targets. Below, we use $\kappa(P)$ to denote the condition number of a matrix $P$.
@@
$\begin{split}\lVert\mathbb{E}\hat{x}_{n+1}\hat{x}_{n+1}^{T}-\alpha_{n+1}\Sigma_{ff}\rVert&\leq\kappa(P_{ff})\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert\mathbb{E}\hat{x}_{1}\hat{x}_{1}^{T}-\alpha_{1}\Sigma_{ff}\rVert+M_{f}\gamma_{n},\\ \lVert\mathbb{E}\hat{y}_{n+1}\hat{y}_{n+1}^{T}-\gamma_{n+1}\Sigma_{ss}\rVert&\leq\kappa(P_{\Delta})\prod_{t=1}^{n}\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert\mathbb{E}\hat{y}_{1}\hat{y}_{1}^{T}-\gamma_{1}\Sigma_{ss}\rVert+\frac{M_{s}}{n}\end{split}$ (12)
@@
The asymptotic distribution of TSA-PR involves the Schur complement $G=A_{ff}-A_{fs}A_{ss}^{-1}A_{ss}$ of $A_{ss}$, as well as the Schur complement $\Delta$ defined in Section 2.
@@
$\begin{split}d_{1}\left(\sqrt{n}G(\bar{x}_{n}-x^{*}),(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right),\\ d_{1}\left(\sqrt{n}\Delta(\bar{y}_{n}-y^{*}),(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right),\end{split}$ (16)
@@
A closed-form solution for $\bar{\Sigma}_{ff}$ and $\bar{\Sigma}_{ss}$ is given in Section C, Eq. (63). The error bound above suggests that an appropriate choice of step sizes is $a=1/2+1/\log n$ and $b=a+1/\log n$ for sufficiently large $n$. With this choice, the error bound is $\mathcal{O}(n^{-1/4})$.
By restricting the test functions $h$ in the Wasserstein-1 distance in Eq. (14) to be $h(x)\in\{-\lVert x\rVert,\lVert x\rVert\}$, we have the following lower and upper bounds on the expected error.
@@
The bounds on the expected error is tight, in the sense that the expected error is bounded above by rate $n^{-1/2}$, and similar to Eq. (13), both $\mathbb{E}\lVert\bar{x}_{n}-x^{*}\rVert$ and $\mathbb{E}\lVert\bar{y}_{n}-y^{*}\rVert$ are lower bounded by the expected norm of the limiting Gaussian vector times the rate $n^{-1/2}$. Replacing the norm of the Gaussian vector with the trace of its covariance would yield $\mathcal{O}(n^{-1/2})$ instead of the $o(n^{-1/2})$ above.
@@ 4 Experiments
The expected errors of $x_{t}$ and $y_{t}$ for every $t\leq n$ are recovered from Theorem 3.1 and Jensen’s inequality; a detailed calculation is shown in Section C.
@@
$\begin{split}\mathbb{E}\lVert G\bar{x}_{n}\rVert-\frac{1}{\sqrt{n}}\mathbb{E}\lVert(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\rVert&\leq\frac{1}{\sqrt{n}}d_{1}\left(\sqrt{n}G\bar{x}_{n},(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right),\\ \mathbb{E}\lVert\Delta\bar{y}_{n}\rVert-\frac{1}{\sqrt{n}}\mathbb{E}\lVert(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\rVert&\leq\frac{1}{\sqrt{n}}d_{1}\left(\sqrt{n}\Delta\bar{y}_{n},(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\right).\end{split}$ (36)
From Theorem 3.2, the upper bound decays faster than $n^{1/2}$, and we have that $\mathbb{E}\lVert G\bar{x}_{n}\rVert$ and $\mathbb{E}\lVert\Delta\bar{y}_{n}\rVert$ are both $\mathcal{O}(n^{-1/2})$. The lower bound is obtained by using the test function $-\lVert x\rVert$.
@@ 6 Conclusion
Acknowledgement
This work was supported by the National Science Foundation under awards ECCS-CAREER-2339509 and CCF-2343599.
@@ A.1 Recursion Updates
Second, we design a sequence of matrices $\{L_{t}\}$ such that the recursion for $\tilde{x}_{t}=\hat{x}_{t}+L_{t}\hat{y}_{t}$ does not depend on $\hat{y}_{t}$ explicitly. In the following, we use $P_{t}$ to denote a coordinate transformation matrix. The choice of $L_{t}$ determines the coordinate transformation
@@ A.2 Usage of Assumptions 13
Here we state a few elementary properties that will be used to prove Theorems 3.1 and 3.2. We will express the difference between finite-time second moments and the asymptotic covariance as a contraction. For a positive definite matrix $P$ and any matrix $A$, we define the weighted norm
$\lVert A\rVert^{2}_{P}=\sup_{x\neq 0}\frac{\langle Ax,Ax\rangle_{P}}{\langle x,x\rangle_{P}}=\sup_{x\neq 0}\frac{x^{T}A^{T}PAx}{x^{T}Px}.$
Note the equivalence of norms
$\lVert A\rVert\leq\sqrt{\kappa(P^{-1})}\lVert A\rVert_{P},\;\lVert A\rVert_{P}\leq\sqrt{\kappa(P)}\lVert A\rVert,$ (42)
where $\kappa(P)$ is the condition number of $P$. For a symmetric matrix $R$, observe
$\lVert R-\alpha(AR+RA^{T})\rVert_{P}$ $=\left\lVert\left(\frac{1}{2}I-\alpha A\right)R+R\left(\frac{1}{2}I-\alpha A^{T}\right)\right\rVert_{P}$
$\leq\lVert I-2\alpha A\rVert_{P}\lVert R\rVert_{P}.$
If $A$ satisfies Assumption 1, i.e., there exists a positive definite matrix $P$ such that
$AP+PA^{T}=I,$
then it is well known (e.g., Lemma 17 (Kaledin et al., 2020)) that for sufficiently small $\alpha$,
$\lVert I-2\alpha A\rVert_{P}\leq\sqrt{1-\frac{1}{2\lVert P\rVert^{2}}\alpha}\leq\left(1-\frac{\mu}{2}\alpha\right),$ (43)
where $\mu=(2\lVert P\rVert)^{-2}$, where the last inequality uses $\sqrt{1-x}\leq 1-x/2$. Hence, we conclude that the following contraction holds:
$\lVert R-\alpha(AR+RA^{T})\rVert_{P}\leq\left(1-\frac{\mu}{2}\right)\lVert R\rVert_{P}.$
@@ A.3 Size of Linear Transformation
A contraction occurs for $L_{t+1}$ in the norm
$\lVert L\rVert_{P_{\Delta},P_{ff}}^{2}=\sup_{x\neq 0}\frac{\langle Lx,Lx\rangle_{P_{ff}}}{\langle x,x\rangle_{P_{\Delta}}},$
where it was shown that
$\lVert L_{t+1}\rVert_{P_{\Delta},P_{ff}}\leq\left(1-\frac{\mu_{ff}}{2}\alpha_{t}\right)\lVert L_{t}\rVert_{P_{\Delta},P_{ff}}+\mathcal{O}(\gamma_{t}).$
By induction and Eq. (44),
$\lVert L_{t+1}\rVert_{P_{\Delta},P_{ff}}$ $\leq\prod_{t=1}^{n}\left(1-\frac{\mu_{ff}}{2}\alpha_{t}\right)\lVert L_{1}\rVert_{P_{\Delta},P_{ff}}+\mathcal{O}\left(\frac{\gamma_{n}}{\alpha_{n}}\right).$
The decay in the unweighted norm holds at the same rate by equivalence of norms.
Appendix B Mean Square Analysis of TSA
@@ B.1 Fast-Time-Scale Mean-Square Convergence
We note that $\Sigma_{ff}$ exists since $-A_{ff}$ is Hürwitz.
@@
$\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert\leq\kappa(P_{ff})\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert\mathbb{E}X_{1}-\alpha_{1}\Sigma_{ff}\rVert+M_{f}\gamma_{n}.$
@@
$\lVert R_{t}^{ff}(X_{t})\rVert_{P_{ff}}\leq M_{f}^{\prime}(\lVert X_{t}\rVert_{P_{ff}}+\alpha_{t})\leq M_{f}^{\prime}\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}+M_{f}^{\prime}\alpha_{t}(1+\lVert\Sigma_{ff}\rVert_{P_{ff}}).$
Such a constant exists because $\lVert\delta_{t}\rVert$ in (48) is uniformly bounded by Proposition A.1. The choice of step sizes ensures Eq. (43) is satisfied for the symmetric matrix $X_{t}-\alpha_{t}\Sigma_{ff}$. By the property $\alpha_{t}-\alpha_{t+1}\leq t^{-1}\alpha_{t+1}$,
$\lVert X_{t+1}-\alpha_{t+1}\Sigma_{ff}\rVert_{P_{ff}}$ $\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{2}+\gamma_{t}M_{f}^{\prime}\right)\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}+\frac{\alpha_{t+1}}{t}\lVert\Sigma_{ff}\rVert_{P_{ff}}$
$+\alpha_{t}\gamma_{t}M_{f}^{\prime}(\lVert\Sigma_{ff}\rVert_{P_{ff}}+1)$
$\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}+\frac{\alpha_{t}}{t}\lVert\Sigma_{ff}\rVert_{P_{ff}}$
$+\alpha_{t}\gamma_{t}M_{f}^{\prime}(\lVert\Sigma_{ff}\rVert_{P_{ff}}+1).$
The preceding equation is of the form
$\lVert X_{t+1}-\alpha_{t+1}\Sigma_{ff}\rVert_{P_{ff}}\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}+\mathcal{O}\left(\alpha_{t}\gamma_{t}+\frac{\alpha_{t}}{t}\right).$ (52)
By induction and Eq. (44), we have that for some constant $M_{f}>0$, it holds that
$\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert_{P_{ff}}$ $\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{1}-\alpha_{1}\Sigma_{ff}\rVert_{P_{ff}}$
$+M_{f}^{\prime\prime}\sum_{t=1}^{n}\alpha_{t}\left(\gamma_{t}+\frac{1}{t}\right)\prod_{j=t+1}^{n}\left(1-\alpha_{j}\frac{\mu_{ff}}{4}\right)$
$\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{1}-\alpha_{1}\Sigma_{ff}\rVert_{P_{ff}}+M_{f}\left(\gamma_{n}+\frac{1}{n}\right)$
Using equivalence of norms, we obtain a bound on the unweighted norm
$\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert$ $\leq\sqrt{\kappa(P_{ff}^{-1})}\lVert X_{n+1}-\alpha_{n+1}\Sigma_{ff}\rVert_{P_{ff}}$
$\leq\kappa(P_{ff})\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert X_{1}-\alpha_{1}\Sigma_{ff}\rVert+\mathcal{O}\left(\gamma_{n}+\frac{1}{n}\right),$
where we re-use the constant $M_{f}$ in the result. By substituting $\hat{x}_{t}=\tilde{x}_{t}-L_{t}\hat{y}_{t}$, we obtain a bound for $\lVert\mathbb{E}\hat{x}_{t}\hat{x}_{t}^{T}-\alpha_{t+1}\Sigma_{ff}\rVert$. A detailed calculation can be found in Section C.